OpenStudy (anonymous):

help please!!! the statement tan theta=-12/5, csc theta=-13/13, and the terminal point determined by theta is in quadrant 3 a. cannot be true because tan theta is greater than zero in quadrant 3 b. cannot be true because if tan theta = -12/5 then csc theta= +- 13/5 c. cannot be true because 12 squared+ 1 squared is not equal to 1 d. cannot be true because tan theta must be less than 1

OpenStudy (anonymous):

\[\tan \theta = \frac{ \sin \theta }{ \cos \theta } \] In quadrant III, are sin/cos positive or negative?

OpenStudy (anonymous):

And quadrant III is the quadrant below the x-axis and to the left of the y-axis. Where all numbers are going to be...

OpenStudy (anonymous):

a. cannot be true because tan theta is greater than zero in quadrant 3